# From time crystals to time quasicrystals: Exploring novel phases in transverse field Ising chains

### v1

@@ V.1 Stability of the TQC phase in the ITF chain
In this section, we aim to systematically investigate the stability of the observed TQC phase within the disordered ITF chain, with a specific focus on the effects of interactions and imperfections, while also assessing the potential role of finite size effects on this TQC phase.
To rigorously confirm that the observed TQC phase in the ITF chain is not merely a finite size effect, it is crucial to assess the stability of this phase as the system size increases. Our exact diagonalization results for chains of lengths 4, 6, 8, and 10 demonstrate a progressive stabilization in the oscillatory behavior of the magnetization amplitude, indicating that larger system sizes lead to more stable oscillations. However, simulating larger chains poses significant computational challenges, particularly when a driving term, like that described by Eq. (9), complicates the dynamics. In this context, utilizing tensor network techniques, such as the time-evolving block decimation (TEBD) algorithm, proves advantageous. This method allows for efficient representation and manipulation of quantum states, making it particularly suited for exploring quantum many-body systems, especially in scenarios where traditional methods falter due to the exponential growth of the Hilbert space.
Furthermore, when the external field $h$ is significantly weaker than the coupling constant $J$, the system enters an MBL phase, characterized by the absence of thermalization and the presence of localized eigenstates. These traits contribute to the stabilization of the quasicrystalline phase as the system size increases. The robustness of the MBL phase against perturbations suggests that a similar resilience may also apply to the TQC phase observed in the ITF chain.
@@ V.2 The stability of the TQC phase under interaction perturbations
We now investigate the stability of the TQC phase in the presence of interaction perturbations. Following a similar approach as before, we incorporate the interacting $\lambda$ term into the ITF Hamiltonian from Eq. (7) as:
@@
The FFT analysis of the magnetization $m^{z}(t)$ is presented in Fig. 8 for different values of the interaction strength $\lambda$. The prominent peaks observed at $k\omega/2\pm\Omega$, where $k$ is an odd integer, suggest that the TQC phase demonstrates resilience to a range of interaction strengths.
Figure 8: Up: FFT analysis of the magnetization $m^{z}$ in the TQC phase for various values of the interaction strength $\lambda$ (up). The data, derived from exact diagonalization of the Hamiltonian for a disordered ITF chain of length 8 with parameters $J=5.5$ and $h=0.3$, are averaged over 200 random realizations. FFTs are computed over a time window of $100T$. Observed peaks at $k\omega/2\pm\Omega$ (where $k$ is an odd integer), demonstrate the resilience of TQC against interaction perturbations. Notably, as $\lambda$ increases, peak amplitudes decrease, ultimately vanishing at a critical value $\lambda_{c}$. Down: Variation of $\Delta A$ across different time windows ($\tau$) for various interaction perturbations. For a chain length of $L=8$, the TQC phase becomes unstable for interactions exceeding $\lambda_{c}=0.03$, as indicated by the observed decay at $\lambda>\lambda_{c}$. Error bars represent the standard deviation of the data, calculated across different time windows.
@@
As the parameter $\lambda$ increases, the peak amplitudes diminish and ultimately vanish at a critical value, $\lambda_{c}$. To determine $\lambda_{c}$, we employ the short-time Fourier transform (STFT) on the magnetization $m^{z}$. We segment the time interval of $100T$ into ten equal windows, performing a Fourier transform within each window to extract the amplitude of the first peak for varying values of $\lambda$. The difference in amplitude of the first peak at a non vanishing $\lambda$ and the amplitude at $\lambda=0$, denoted as $\Delta A$, provides valuable insight into the critical value $\lambda_{c}$. As shown in Fig. 8-(down), $\Delta A$ approaches zero for $\lambda$ values below $\lambda_{c}$, indicating that the TQC remains stable against interactions weaker than $\lambda_{c}$. For a chain of length $L=8$, the critical interaction value is approximately $0.03$, as illustrated in Fig. 8-(down).
The TQC phase is not merely a finite-size effect; it manifests in both finite disordered ITF chains and infinite chains alike. To assess the stability against interactions, we examined a chain of length $L=8$. It is evident that since the system resides in the MBL phase for $h<J$, we anticipate that the stability of the TQC phase will improve with increasing system size.
@@ V.3 The stability of the TQC phase under imperfect rotations
where $\epsilon$ cause imperfect rotation of spins. The modified drive term introduces a scaling factor $(1-\epsilon)$ to the $\cos(\Omega t)$ term, which captures the effect of imperfect rotations. The FFT of the magnetization $m^{z}$ is illustrated in Fig.9 for various values of the parameter $\epsilon$. The impact of imperfection is similar to the interaction perturbation. A key observation from the FFT of the magnetization $m^{z}$ presented in Fig. 9 is that the presence of imperfections leads to a reduction in the amplitude of the subharmonic peaks without affecting their positions. Specifically, these subharmonic peaks appear at locations given by $k\frac{\omega}{2}\pm\Omega$ where $k$ is an odd integer, indicating that while the system’s frequency structure remains stable, the strength of the response diminishes as $\epsilon$ increases.
Figure 9: Up: FFT analysis of the magnetization $m^{z}$ in the TQC phase for various values of the imperfection $\epsilon$. The data, derived from exact diagonalization of the Hamiltonian for a disordered ITF chain of length 8 with parameters $J=5.5$ and $h=0.3$, are averaged over 200 random realizations. FFTs are computed over a time window of $100T$. Observed peaks at $k\omega/2\pm\Omega$ (where $k$ is an odd integer), demonstrating the resilience of TQC against imperfections. Notably, as $\epsilon$ increases, peak amplitudes decrease, ultimately vanishing at a critical value $\epsilon_{c}$ ($\sim 0.01$ for a chain of length 8). Down: Variation of $\Delta A$ across different time windows ($\tau$) for various imperfections. For a chain length of 8, the TQC phase becomes unstable for imperfections exceeding $\epsilon_{c}=0.01$, as indicated by the observed decay at $\epsilon>\epsilon_{c}$.
@@
To determine $\epsilon_{c}$, we study the behavior of the difference in the amplitude of the first peak at a non vanishing $\epsilon$ and the amplitude at $\epsilon=0$, denoted as $\Delta A=A(\epsilon\neq 0)-A(\epsilon=0)$, for a chain of length 8. As shown in Fig. 9-(down), $\Delta A$ approaches zero for $\epsilon$s below $\epsilon_{c}=0.01$, indicating that the TQC remains stable against imperfections smaller than $0.01$.
@@
As demonstrated, the critical perturbation thresholds for stability in TCs are significantly higher, by an order of magnitude, than those observed in TQCs, indicating that TCs exhibit markedly greater robustness to perturbations. This enhanced robustness is attributed to the phenomenon of quantum frequency lockingr27 , which occurs when a quantum system with an inherent characteristic frequency $\omega_{c}$ is driven at a frequency $\Omega=2\pi/T$ that is close to a rational multiple $q/r$ of $\omega_{c}$. Under these conditions, the system can achieve stable oscillations with a period precisely aligned to an integer multiple of the driving period, $qT$. In contrast, TQCs do not benefit from this frequency locking, leading to magnetization characterized by incommensurate oscillations.
VI Absence of TQC phase in ITF chain with $J_{i}$ in symmetric intervals
@@
In this section, we investigate the absence of the TQC phase in the ITF model described by Eq. (7), where the exchange couplings $J_{i}$ are selected from a symmetric interval of $[-J/2,J/2]$. This selection implies that the exchange interactions can take both positive and negative values, leading to a rich landscape of magnetic behavior.
@@
Despite the absence of net magnetization, the system is not in a trivial paramagnetic phase; instead, it resides in a spin glass (SG) phaser24 ; r25 ; r26 . This conclusion is supported by the observation that the Edwards-Anderson order parameter is non-zero EA , indicating the presence of localized magnetic order amidst the disordered couplings. In a spin glass phase, the system can sustain a finite density of domain walls, which become immobilized due to the underlying disorder. This immobilization results in locally broken symmetry, as different regions of the system may adopt distinct magnetic configurationsr23 ; D.Huse .
As a result, the eigenstates of the Hamiltonian are characterized by random glassy configurations, showcasing a complex interplay between disorder and magnetic interactions. These configurations reflect the frustration and variability inherent in the system, leading to a diverse array of possible states that can be accessed under different conditions. This behavior is characteristic of spin glasses, where the combination of randomness and interaction leads to a non-trivial energy landscape, challenging conventional notions of order in magnetic systems.
@@
where, each spin variable $\sigma_{i}$ can randomly take values of $+1$ or $-1$, with the constraint that $\bar{\sigma}=-\sigma$. Under the influence of magnetic fields, the system experiences tunneling between the states $|j_{1},j_{2}\rangle$ and $|\bar{j}_{1},\bar{j}_{2}\rangle$. As a result, the Hamiltonian’s eigenstate emerges as a coherent superposition of these states as:
@@
The driven ITF model exhibits rich phases, particularly when compared to its undriven counterpart. In the undriven disordered ITF model, the primary phases observed are the PM and SG phases. However, the introduction of driving field leads to the emergence of additional dynamical phases.
@@
1. $\pi$PM-MBL phase: This phase is characterized by a paramagnetic behavior under the influence of a periodic driving field, where the system exhibits localization properties typical of MBL phases. In this regime, the quantum coherence and entanglement are preserved, despite the presence of disorder and external driving.
2. $\pi$SG-MBL phase: In this phase, the system displays spin glass characteristics while still retaining MBL traits. The dynamics are influenced by the interplay between the disorder, the transverse field, and the driving, leading to a scenario where the system can remain localized while exhibiting complex spin configurations typical of spin glass states.
In the periodically driven spin glass phase, our findings (not shown) demonstrate the stability of the TC phase even amidst interaction perturbations and imperfect rotations. The system persists in the localized phase for interaction strengths ranging from $0<\lambda\leq 1$ r7 , highlighting its resilience to these perturbations while maintaining its position in the TC phase.
@@
Our investigations into the time evolution of the magnetizations $m^{x,y,z}$ indicate a notable absence of TQC in either the SG-MBL or the PM-MBL phases under the current configuration. This observation is in stark contrast to scenarios where $J_{i}$ is selected from the range $[J/2,3J/2]$, where quasiperiodic dynamics can manifest. The lack of a TQC in the present setup can be primarily attributed to the characteristics inherent to the spin glass phase. The presence of domain walls disrupts the potential for the system to exhibit dynamics that repeat in a quasiperiodic fashion over time. Consequently, there are no sustained oscillations in magnetization with a robust frequency, underscoring the intricate interplay between disorder and driving in shaping the dynamical properties of spin systems.
This highlights the complexity of quantum systems under external influences and the significant role that disorder plays in determining their dynamical behavior, ultimately uncovering a complex interplay of phase interactions that deserves deeper investigation.
VII Conclusion
In summary, our investigation provides a thorough understanding of the dynamical properties of the disordered Ising model in a transverse magnetic field (ITF) under periodic and quasiperiodic driving conditions.
In the first part of our study, we demonstrated that the driven disordered ITF chain exhibits time-dependent magnetization $m^{z}$ characterized by significant oscillations with a period of $2T$ and an oscillatory-decaying envelope. This stability against various perturbations highlights the robustness of the time crystal (TC) phase, indicating its potential applications in quantum technologies.
In the second part, we explored the ability of the disordered ITF chain to realize a time quasicrystalline (TQC) phase. By applying a sinusoidal driving field with frequencies that are irrationally related, we observed complex frequency structures in the magnetization that resemble those of quasicrystals. This finding paves the way for further exploration of TQC in disordered systems, showcasing the rich dynamical behavior that can arise from quasiperiodic driving.
Our work establishes a foundation for future experimental implementations in platforms such as cold atomic systems and quantum simulators. The insights obtained from TQCs hold significant implications for quantum information, quantum computing, and the development of novel materials with specific dynamical responses. The interplay between disorder, driving, and quantum coherence in these systems calls for further investigation, potentially leading to breakthroughs in understanding and harnessing complex quantum behaviors.

### v2

@@ V.1 Stability of the TQC phase in the ITF chain
In this section, we will analyze the stability of the TQC phase when subjected to different perturbations. Prior to this, it is helpful to review how a non-equilibrium system thermalizes under various types of driving fields.
Local time-dependent Hamiltonians can only induce local rearrangements. In strongly disordered systems, such rearrangements require a nonzero energy cost and are typically nonresonant with the discrete harmonics of the periodic drive, thereby preserving MBL. In contrast, random driving, with its continuous frequency spectrum, can resonantly activate arbitrary local transitions, leading to energy absorption and eventual thermalization. Consequently, MBL is generally unstable under random driving. Quasiperiodic drives, which consist of multiple incommensurate discrete frequencies, create a noncontinuous but dense spectrum that can preserve MBL under typical conditions. However, over long times, they may lead to slow thermalization of the system. However, over long timescales, which diverge exponentially for weak or rapid drives, the system eventually thermalizes to infinite temperature. Nonetheless, slow relaxation processes can give rise to metastable dynamical phases, reminiscent of prethermal regimes dumitrescu2018 ; zhao2021 .
The presence of a long-lived quasi-MBL regime with only slow decay suggests the emergence of transient phases unique to quasiperiodically driven systems. Analogous to prethermal state, this regime can be interpreted as a prethermal TQC phase. While the system does eventually thermalize, the process is sufficiently slow to allow the emergence of a TQC phase. The prethermalization lifetime depends on several parameters, notably the drive frequency, where higher frequencies prolong the prethermal regime. Importantly, under quasi-periodic driving, the heating dynamics are highly sensitive to the drive waveform smoothness; rectangular pulses markedly increase energy absorption and accelerate thermalization, whereas smoother waveforms such as sinusoidal drives substantially enhance the stability and duration of the prethermal state. Additionally, the drive amplitude critically influences the system’s energy uptake, with larger amplitudes expediting thermalization rr17 ; dumitrescu2018 ; zhao2021 .
Following, we systematically investigate the stability of the observed TQC phase within the disordered ITF chain, with a specific focus on the effects of interactions and imperfections.
@@ V.2 The stability of the TQC phase under interaction perturbations
In order to investigate the stability of the TQC phase in the presence of interaction perturbations, following a similar approach as before, we incorporate the interacting $\lambda$ term into the ITF Hamiltonian from Eq. (8) as:
@@
The FFT analysis of the magnetization $m^{x}(t)$ is presented in Fig. 7 for different values of the interaction strength $\lambda$. The prominent peaks observed at $k\omega/2\pm\Omega$, where $k$ is an odd integer, suggest that the TQC phase demonstrates resilience to a range of interaction strengths.
Figure 7: Up: FFT analysis of the magnetization $m^{x}(t)$ in the TQC phase for various values of the interaction strength $\lambda$ (up). The data, derived from exact diagonalization of the Hamiltonian for a disordered ITF chain of length 8 with parameters $J=5.5$ and $h=0.3$, are averaged over 200 random realizations. FFTs are computed over a time window of $100T$. Observed peaks at $k\omega/2\pm\Omega$ (where $k$ is an odd integer), demonstrate the resilience of TQC against interaction perturbations. Notably, as $\lambda$ increases, peak amplitudes decrease, ultimately vanishing at a critical value $\lambda_{c}$. Down: Variation of $\Delta A$ across different time windows ($\tau$) for various interaction perturbations. For a chain length of $L=8$, the TQC phase becomes unstable for interactions exceeding $\lambda_{c}=0.03$, as indicated by the observed decay at $\lambda>\lambda_{c}$. Error bars represent the standard deviation of the data, calculated across different time windows.
@@
As the parameter $\lambda$ increases, the peak amplitudes diminish and ultimately vanish at a critical value, $\lambda_{c}$. To determine $\lambda_{c}$, we employ the short-time Fourier transform on the magnetization $m^{x}(t)$. We segment the time interval of $100T$ into ten equal windows, performing a Fourier transform within each window to extract the amplitude of the first peak for varying values of $\lambda$. The difference in amplitude of the first peak at a non vanishing $\lambda$ and the amplitude at $\lambda=0$, denoted as $\Delta A$, provides valuable insight into the critical value $\lambda_{c}$. As shown in Fig. 7-(down), $\Delta A$ approaches zero for $\lambda$ values below $\lambda_{c}$, indicating that the TQC remains stable against interactions weaker than $\lambda_{c}$. For a chain of length $L=8$, the critical interaction value is approximately $0.03$, as illustrated in Fig. 7-(down).
@@ V.3 The stability of the TQC phase under imperfect rotations
where the parameter $\epsilon$ quantifies the degree of imperfection in the spin rotations. The inclusion of the factor $(1-\epsilon)$ in front of the $\cos(\Omega t)$ term models the deviation from ideal rotation, effectively capturing the impact of imperfect control on the system’s dynamics.
The FFT of the magnetization $m^{x}(t)$ is illustrated in Fig. 8 for various values of the parameter $\epsilon$. The impact of imperfection is similar to the interaction perturbation. A key observation from the FFT of the magnetization $m^{x}(t)$ presented in Fig. 8 is that the presence of imperfections leads to a reduction in the amplitude of the subharmonic peaks without affecting their positions. Specifically, these subharmonic peaks appear at locations given by $k\frac{\omega}{2}\pm\Omega$ where $k$ is an odd integer, indicating that while the system’s frequency structure remains stable, the strength of the response diminishes as $\epsilon$ increases.
Figure 8: Up: FFT analysis of the magnetization $m^{x}(t)$ in the TQC phase for various values of the imperfection $\epsilon$. The data, derived from exact diagonalization of the Hamiltonian for a disordered ITF chain of length 8 with parameters $J=5.5$ and $h=0.3$, are averaged over 200 random realizations. FFTs are computed over a time window of $100T$. Observed peaks at $k\omega/2\pm\Omega$ (where $k$ is an odd integer), demonstrating the resilience of TQC against imperfections. Notably, as $\epsilon$ increases, peak amplitudes decrease, ultimately vanishing at a critical value $\epsilon_{c}$ ($\sim 0.01$ for a chain of length 8). Down: Variation of $\Delta A$ across different time windows ($\tau$) for various imperfections. For a chain length of 8, the TQC phase becomes unstable for imperfections exceeding $\epsilon_{c}=0.01$, as indicated by the observed decay at $\epsilon>\epsilon_{c}$.
@@
[…]
@@
The critical perturbation thresholds for stability in TCs are significantly higher, by an order of magnitude, than those observed in TQCs, indicating that TCs exhibit markedly greater robustness to perturbations. This enhanced robustness is attributed to the phenomenon of quantum frequency locking r27 , which occurs when a quantum system with an inherent characteristic frequency $\omega_{c}$ is driven at a frequency $\Omega=2\pi/T$ that is close to a rational multiple $q/r$ of $\omega_{c}$. Under these conditions, the system can achieve stable oscillations with a period precisely aligned to an integer multiple of the driving period, $qT$.
In contrast, frequency locking in quasiperiodically driven systems does not occur in the simple and regular manner observed under periodic driving. Instead, more intricate forms of synchronization can arise when the driving frequencies are incommensurate (i.e., their ratios are irrational numbers). An analogous effect, sometimes referred to as irrational mode locking, has been reported in systems with a quasiperiodic spatial background pre23 . While the physical setting there differs from the temporally quasiperiodic driving considered in our study, our results demonstrate that a related mechanism can manifest in time, leading to richer frequency spectra. In particular, quasiperiodic driving allows the system’s response to lock onto linear combinations of the driving frequencies with integer coefficients but irrational overall ratios. This produces quasi stable oscillatory behavior that is structured yet aperiodic, reflecting locking onto these irrational frequency combinations.
V.4 The prethermal TQC lifetime: effects of different system-sizes
To examine the effect of system size on the lifetime of the prethermal TQC in the ITF chain, we compare the time evolution of the magnetization for chains of length $L=6,8$, and 10, as illustrated in Fig. 9. The results show that the TQC lifetime is approximately independent of the system size over this range.
As shown in Ref. Kemp2017 , the edge spin in finite-size, open-boundary integrable spin chains—such as the ITF chain and the spin-$1/2$ XYZXYZ chain—can preserve the memory of its initial state over remarkably long times. Specifically, the expectation value of $\sigma^{z}_{1}$, representing the edge spin, stabilizes after an initial transient into a plateau that defines the edge spin coherence time. This coherence time, which originates from strong zero modes, grows exponentially with system size and depends sensitively on parameters like the exchange interaction $J$ and the transverse field $h$.
In our quasiperiodically driven model, the magnetization envelope also forms a plateau after an initial transient, characterizing the prethermal TQC lifetime. This plateau plays a role analogous to the edge coherence time described in Kemp2017 . However, unlike the integrable case, the duration of this plateau does not show a pronounced dependence on system size, at least for the sizes $L=6,8,10$ we studied. Instead, its lifetime is primarily controlled by drive parameters, such as the structure of the quasiperiodic modulation (see Fig. 10), which dominate the stability of the prethermal TQC regime.
Figure 9: The magnetization $m^{x}(t)$ for different system sizes at $\omega=1$ and $\Omega=\sqrt{2}/4$. The data show that the prethermal TQC lifetime, defined as $t_{\mathrm{th}}=\tau^{*}-\tau_{\mathrm{pre}}$, is approximately independent of $L$ within our numerical accuracy.
V.5 The prethermal TQC lifetime: effects of derive frequencies
In this section, we systematically examine how the prethermal time $t_{th}$ depends on the drive frequencies, by varying $\omega$ while maintaining a constant ratio $\Omega/\omega=\sqrt{2}/4$. This approach allows us to isolate the influence of an overall frequency scale without modifying the relative structure of the quasiperiodic driving.
As shown in Fig. 10, increasing $\omega$ extends the plateau region between $\tau_{\mathrm{pre}}$ and $\tau^{*}$, which prolongs the prethermal TQC lifetime by delaying its thermalization.
Figure 10: Time evolution of the magnetization $m^{x}(t)$ for a quasiperiodically driven ITF chain of length $L=6$, with exchange couplings $J_{i}$ sampled from $[J/2,3J/2]$ ($J=5.5$), and a transverse field $h=0.3$. The ratio is fixed at $\Omega/\omega=\sqrt{2}/4$. Panels (a) to (d) correspond to driving frequencies of $\omega=1.0$, $1.25$, $1.5$, and $1.75$, respectively. The TQC lifetime increases with the driving frequency $\omega$, indicating that higher drive frequencies suppress heating, thereby enhancing the prethermal TQC lifetime.
The extension of the prethermal TQC lifetime is primarily due to the suppression of energy absorption processes in the high-frequency ($\omega$) limit, a characteristic feature of prethermalization in driven many-body systems. Specifically, increasing the driving frequency diminishes the probability of resonant transitions that could otherwise lead to rapid energy delocalization and heating. As a result, the system remains confined within a dynamically constrained manifold for an extended period, exhibiting robust quasiperiodic oscillations despite the absence of exact periodicity in the driving.
The extension of the prethermal TQC lifetime is more pronounced when $\omega\gg J$, where $J$ is the system’s characteristic energy scale. This underscores the importance of a clear separation of timescales $\omega$ and $J$ ($\hbar=1$) for stabilizing the TQC phase. Our findings align with established theoretical arguments for prethermal behavior under high-frequency driving, applicable to both periodically and quasiperiodically driven systemspre1 ; pre2 ; pre3 ; pre4 .
The current results underscore the crucial influence of the driving frequencies in determining the lifetime of the prethermal TQC. These findings not only reaffirm the prethermal character of the observed TQC dynamics but also provide a practical approach to enhance their experimental detectability through tuning the drive frequency.
VI TQC phase in the ITF model with symmetric sampling of $J_{i}$s
@@
In this section, we examine the emergence of the TQC phase within the ITF model outlined in Eq. (8), where the exchange couplings $J_{i}$ are randomly chosen from the symmetric interval [J/2,J/2][-J/2,J/2]. This symmetric sampling implies that the exchange interactions can take both positive and negative values, leading to a rich landscape of magnetic behavior.
@@
Despite the absence of net magnetization, the system is not in a trivial paramagnetic phase; instead, it resides in a spin glass (SG) phase r24 ; r25 ; r26 . This conclusion is supported by the observation that the Edwards-Anderson order parameter is non-zero EA , indicating the presence of localized magnetic order.
In a spin glass phase, the system can sustain a finite density of domain walls, which become immobilized due to the underlying disorder. This immobilization results in locally broken symmetry, as different regions of the system may adopt distinct magnetic configurations r23 ; D.Huse . As a result, the eigenstates of the Hamiltonian are characterized by random glassy configurations, showcasing a complex interplay between disorder and magnetic interactions. These configurations reflect the frustration and variability inherent in the system, leading to a diverse array of possible states that can be accessed under different conditions. This behavior is characteristic of spin glasses, where the combination of randomness and interaction leads to a non-trivial energy landscape, challenging conventional notions of order in magnetic systems.
@@
where, $\sigma_{i}$ can randomly take values of +1+1 or 1-1, with the constraint that $\bar{\sigma}=-\sigma$. The above states are the eigenstate of the Ising Hamiltonian $H=\sum_{i}J_{i}\sigma_{i}\sigma_{i+1}$, with random $J_{i}$. Under the influence of magnetic fields, the system experiences tunneling between the states $|j_{1},j_{2}\rangle$ and $|\bar{j}_{1},\bar{j}_{2}\rangle$. As a result, the Hamiltonian’s eigenstate emerges as a coherent superposition of these states as:
@@
The driven disordered ITF model exhibits rich phases, particularly when compared to its undriven counterpart. In the undriven disordered ITF model, the primary phases observed are the PM and SG phases. However, the introduction of driving field leads to the emergence of additional dynamical phases.
@@
[…]
In the periodically driven spin glass phase, our findings demonstrate a TC phase despite interaction perturbations and imperfect rotations. The system persists in the localized phase for interaction strengths ranging from $0<\lambda\ll 1$ r7 , highlighting its resilience to these perturbations while maintaining its position in the TC phase.
Shifting focus to the quasiperiodic spin chain governed by the Hamiltonian in Eq. (8), we examine the time-dependent Hamiltonian $\hat{H}_{ITF}(t)$ as expressed in Eq. (9). The couplings $J_{i}$ are drawn from the range [J/2,J/2][-J/2,J/2], and the external fields $h_{i}$ are selected from $[0,h]$. The effects of the external driving field, represented by the Hamiltonian $\hat{H}_{drive}(t)$, are also incorporated into the analysis.
The time evolution of the magnetization, as depicted in Fig. 11 (up), reveals a notable absence of TQC in either the SG or the PM phases under the condition of $\omega=1$ and $\Omega=\sqrt{2}/4$. This observation is in stark contrast to scenarios where $J_{i}$ is selected from the range $[J/2,3J/2]$, where quasiperiodic dynamics can manifest.
Figure 11: Time evolution of the magnetization $m^{x}(t)$ for a quasiperiodically driven ITF chain of length $L=8$. The exchange couplings $J_{i}$ are sampled symmetrically from [J/2,J/2][-J/2,J/2], with $J=5.5$, and the transverse field is set to $h=0.3$. (Up) $\omega=1$; (down) $\omega=12$, with $\Omega/\omega=\sqrt{2}/4$ in both cases.
The lack of a TQC in the present setup can be primarily attributed to the intrinsic properties of the spin glass phase. Specifically, when $\delta J$ is drawn from a symmetric distribution, fluctuations in the couplings equally promote and inhibit local interactions, enhancing energy delocalization and accelerating thermalization.
Interestingly, we find that increasing the driving frequencies $\omega$ and $\Omega$, while maintaining a fixed ratio $\Omega/\omega=\sqrt{2}/4$, leads to more pronounced and longer-lived quasiperiodic magnetization oscillations (see Fig.11-down). This observation highlights that both the choice of the coupling distribution $J_{i}$ and the values of the driving frequencies play a crucial role in stabilizing the TQC phase.
The dynamical robustness of the TQC phase is highly sensitive to the drive structure and the underlying disorder. A detailed understanding of the stability boundaries and the influence of drive parameters, on the lifetime of the prethermal TQC remains an open and promising avenue for future research.
VII Conclusion and outlook
In summary, our investigation provides a detailed understanding of the dynamical response of a disordered quantum Ising chain in a transverse magnetic field (ITF) under both periodic and quasiperiodic driving. In the periodically driven scenario, we confirmed that the system exhibits a robust discrete time crystal (TC) phase, characterized by persistent period-doubled oscillations in the magnetization $m^{z}(t)$. These oscillations remain stable despite the presence of interaction perturbations and imperfections in the driving protocol. Additionally, we demonstrated that increasing the chain length further enhances the stability of the TC phase.
In the quasiperiodically driven case, we observe the emergence of complex, structured oscillations at multiple incommensurate frequencies. These oscillations persist for a long time but eventually decay, indicating that the TQC phase corresponds to a transient prethermal dynamical regime rather than a stable thermodynamic phase.
The prethermal TQC lifetime is determined by the duration of the first plateau in the magnetization envelope. Our finite-size analysis reveals that this lifetime is largely independent of system size, at least for the chain lengths studied, demonstrating the robustness of the prethermal quasiperiodic order against finite-size effects.
Furthermore, the prethermal TQC lifetime exhibits a strong dependence on the quasiperiodic drive frequency. Our results show that increasing the drive frequency significantly extends the TQC lifetime.
Moreover, we found that TQC phase is suppressed when the exchange couplings are sampled symmetrically around zero, pointing to the destructive role of domain walls and emphasizing the importance of disorder structure. Notably, the quasiperiodic response remains stable under interaction perturbations and rotation imperfections, indicating the robustness of the observed prethermal regime over experimentally relevant timescales. Interestingly, we also observe that increasing the drive frequencies can lead to the reemergence of magnetization oscillations, suggesting that the drive frequencies play a pivotal role in enhancing the lifetime of the TQC phase. This behavior highlights the interplay between the drive parameters and disorder in shaping the dynamical stability of TQCs.
Our results enrich the current understanding of TC behavior in interacting spin systems and suggest clear signatures for experimental realization in platforms such as cold atomic gases and quantum simulators. The dynamical interplay between disorder, driving, and coherence revealed in our work opens avenues for studying novel non-equilibrium phenomena.
An intriguing future direction, inspired by time-quasiperiodic topological superconductors pre18 , is to investigate whether TQC phases in spin chains can host robust edge modes or pairing operators. While such features have been predicted in fermionic systems via mechanisms like Majorana multiplexing, their analogs in spin systems with quasiperiodic driving remain an exciting open question for future exploration.
While a systematic determination of heating rates under general quasiperiodic drives remains an open problem, our results suggest that for the specific driving protocol considered here, such a characterization could be feasible and is an interesting direction for future work, particularly to better understand the persistence and decay of quasiperiodic oscillations in relation to energy absorption. In particular, identifying the precise boundaries of drive parameters that critically influence the lifetime of the prethermal TQC remains an open question.
Quasiperiodicity combined with topology can give rise to robust edge modes; notably, Majorana modes, topologically protected excitations, are central to realizing exotic and stable quantum phases pre18 ; pre19 . In the context of the disordered ITF chain studied here, we have not identified definitive evidence of well-defined edge states or pairing operators associated with the TQC phase. Nonetheless, the question of whether dynamically protected edge states can exist in quasiperiodically driven interacting spin systems remains highly intriguing.
For instance, in the paper pre19 , time-quasiperiodic topological superconductors were investigated, demonstrating that a system driven at $d$ mutually irrational frequencies can host up to $2^{d}$ types of Majorana modes. These modes can coexist despite spatial overlap and the absence of time-translational invariance. The authors introduced the concept of Majorana multiplexing, whereby quasiperiodic driving induces multiple Majorana edge modes at different quasienergies related to the incommensurate drive frequencies. This indicates that quasiperiodicity does not necessarily hinder the formation of topologically protected edge structures.
Moreover, it has been shown that Majorana edge modes, both zero-energy and those at golden-ratio quasienergy, can also appear in spin systems, such as the Fibonacci-driven quantum Ising chain pre18 .
However, extending these concepts to disordered, many-body interacting spin systems under quasiperiodic driving with multiple irrational frequencies remains an open challenge. In particular, whether an analogue of Majorana multiplexing or other topological edge phenomena can manifest in many-body localized or prethermal spin chains driven quasiperiodically is still unknown. This question is particularly relevant to our model, where a disordered interacting Ising chain is subjected to a quasiperiodic drive involving multiple incommensurate frequencies.